OpenStudy (anonymous):

Help Please! Subject: Hypotenuse Q. 3 of 5 @ali2x2 @ab864

OpenStudy (anonymous):

Question?

OpenStudy (anonymous):

18^2 - 4^2 = ? squrt(?) = answer

OpenStudy (anonymous):

324 - 16 = b^2 308 = b^2 17.5499.... = b